# Lessons learned from field demonstrations of model predictive control and reinforcement learning for residential and commercial HVAC: A review

### v1

@@ Appendix A Field demonstration summary tables
MPC Sargans, Switzerland three of four heating circuits in a medium-sized office building hydronic HVAC system with oil-fired boiler providing water-based heating 46 minimize energy 31.9% energy vs. measured baseline
@@
MPC Bordeaux, France open-plan conference room in a University of Bordeaux building radiant underfloor heating, AHU, and FCU for airflow and auxiliary heating 30 minimize energy 30–40% energy vs. simulated baseline
@@
MPC Long Beach, CA; Dayton, OH; NYC; Washington, DC office, courthouse, hospital, and high school buildings forced-air systems with variable frequency drive supply fans 108–170 minimize energy -0.4–8.9% energy vs. measured baseline
@@
MPC West Lafayette, IN, US open-plan office space in Ray W. Herrick Laboratories, Purdue University hydronic radiant floor system, air-cooled chiller assumed for generalization 11 minimize energy 27% energy vs. simulated baseline
@@
MPC West Lafayette, IN, US and FL, US a conference room at Ray W. Herrick Laboratories at Purdue University and a small retail store in Florida forced-air systems with packaged rooftop units in a laboratory conference room and a small retail store 3 and 60 minimize energy and peak demand 14% & 12.95% energy and 50% & 14.45% peak demand vs. measured baseline
@@
MPC West Lafayette, IN, US open-plan office spaces in Ray W. Herrick Laboratories, Purdue University hydronic radiant floor system for the first office, forced air delivery systems for the other two 10 minimize energy and cost 7.81–58% energy and 34.72–75% cost vs. measured and simulated baseline
@@
MPC West Lafayette, IN, US open-plan office space in Ray W. Herrick Laboratories, Purdue University hydronic radiant floor cooling, AHU ventilation, air-cooled chiller assumed for generalization 19 minimize thermal dissatisfaction and energy consumption reduced thermal dissatisfaction with 23.13% energy increase vs. measured baseline
@@
RL Pittsburgh, PA, US The Intelligent Workplace (IW), a single-level office building radiant window system integrating hot water pipes with mullions, forming a novel water-based heating system 78 minimize energy 26.7% energy vs. measured baseline
@@
MPC Singapore experimental cells in the Building and Construction Authority SkyLab hybrid forced air and radiant delivery through active chilled beam systems; primary AHU with induced air mixing 1 minimize energy 14.7–20% energy vs. measured baseline
@@
MPC Singapore single lecture room in the North Academic Complex, Nanyang Technological University (NTU) forced air VAV system with central chiller, DOAS, modified coil, pump, and valves 14 minimize energy 16–20% energy vs. measured baseline
@@
MPC Varennes, QC, Canada The CanmetENERGY Centre, a single-story building with workstations and meeting rooms forced air system with electric boiler for heating and natural gas boilers as backup 70 minimize energy 22.2% energy vs. measured baseline
@@
MPC Hasselt, Belgium three office floors of the HybridGEOTABS office building (Hollandsch Huys) radiant and forced air system with TABS, floor heating, AHUs, and a ground-source heat pump 35 minimize energy 53.5% energy savings vs. measured baseline
@@
MPC Singapore office in Research Techno Plaza (RTP) and a lecture theater, both at NTU forced air system with a ceiling-mounted VAV box and an external FCU 10 minimize energy 36.7–58.5% energy savings vs. measured baseline
@@
RL Pittsburgh, PA, US a conference room at Carnegie Mellon University (CMU) forced air VAV system with variable-speed cooling airflow 21 minimize energy 16.7% energy savings vs. measured baseline
@@
MPC Hamburg, Germany an office building serving as headquarters for two local ministries radiant heating with hot water supplied from a district network 90 minimize energy 30% energy savings vs. measured baseline
@@
MPC Brussels, Belgium five-floor building forced air system with gas boilers, refrigeration units, and AHUs 10–17 minimize energy 8.6–11.7% energy savings vs. measured baseline
@@
MPC Singapore experimental cells in the Building and Construction Authority SkyLab hybrid forced air and radiant cooling via active chilled beams. no heating. Electrochromic window for heat gain control. 1–2 minimize energy 18.4–23.8% energy savings vs. measured baseline
@@
RL Berkeley, CA, US office room at FLEXLAB in LBNL forced air system with variable-capacity AHU, air mixing dampers, shared chiller, PV, and battery 7 minimize cost 0.6–39.6% cost vs. measured baseline
@@
RL Seoul, South Korea climate chamber at Yonsei University $<1$ increase comfort 10.9% reduction in thermal discomfort (PPD) vs. measured baseline
@@
MPC Heverlee, Belgium a small, unoccupied test module in a research facility called the Vliet building 57–69 compare MPC modeling approaches no single model was best
@@
MPC West Lafayette, IN, US a conference room at Ray W. Herrick Laboratories, Purdue University forced air system with single-speed rooftop units and diffusers 5–15 minimize cost 3.61–8.67% cost savings vs. measured baseline
@@
MPC Berkeley, CA, US two floors of Building 59 on the main campus of LBNL, closed offices on the ground floor, open offices on the second forced air underfloor system with DX coils, scroll compressors, cooling towers, and heating coils 31 minimize energy 40% energy savings vs. measured baseline
@@
RL Nashville, TN, US three-story, LEED Gold Certified building with offices, classrooms, halls, a gym, a lounge, and a cafeteria forced air system with a dedicated outside air unit, variable refrigerant flow, and simultaneous AHUs 365 minimize energy 13.1–14.3% energy savings vs. measured baseline
@@
MPC Merced, CA, US central cooling plant of the University of California, Merced forced air system, chilled water, thermal storage, variable-speed pumps, PV, flexible chillers 7 maximize solar self-consumption and minimize peak demands 26.4% increase in solar self-consumption and 9.8% reduction in peak demand vs. measured baseline
@@
MPC Blue Lake Rancheria, CA, US a convenience store with an associated gas station forced air system via two RTUs, refrigeration, freezer systems, and integration with battery and PV 365 minimize cost and peak demands 11.7% cost savings and 34% reduction in peak demand vs. measured baseline
@@
MPC Campo Grande, Brazil operating and waiting rooms in ophthalmology section of a hospital forced air system with AHUs, fixed-speed fans, and variable-speed chiller 4 minimize energy experimental savings unreported; simulated savings 4.76% vs. simulated baseline
@@
RL Singapore an office space in SDE4 building at National University of Singapore (NUS) hybrid system with DOAS and ceiling fans, featuring variable speed control and reduced airflow 10 minimize energy 14% energy savings vs. measured baseline
@@
RL US university campus buildings and a multi-purpose building with shops, restaurants, apartments, and a clinic hybrid system with chiller plant, air handlers, fan coils, pumps, cooling towers, and ductwork 90 minimize energy 9–13% energy savings vs. measured baseline
@@
MPC Singapore six office zones on level 5 of a tropical net-zero energy building forced air system with two FCUs and zone-specific VAVs 49 improve self utilization of PV power and building self-sufficiency 19.5% improvement in PV self-utilization and 10.6% increase in building self-sufficiency vs. simulated baseline
@@
MPC West Lafayette, IN, US private office spaces in Ray W. Herrick Laboratories, Purdue University forced air system, individual VAV boxes, central AHU with reheat coils 1 minimize energy 28–35% energy savings vs. measured baseline
@@
MPC Zhejiang province, China three-floor airport terminal forced air system. air-cooled heat pumps, chilled water pumps, and AHUs 15 minimize energy 24% energy savings vs. measured baseline
@@
MPC Costa Mesa, CA, US xix classrooms in two buildings in a K-12 school forced air system with six rooftop units serving six classrooms 30 minimize peak demand 22–24% total peak reduction vs. measured baseline
@@
MPC Lausanne, Switzerland a multipurpose academic building, the Polydome, on the EPFL campus forced air heat pump system 4 minimize energy 18.4% energy savings vs. measured baseline
@@
MPC ON and QC, Canada central heating plants radiant steam-based system. non-condensing gas boilers for heating and hot water. 48–67 minimize energy, cost, and GHG emissions 1.3–2.8% energy and cost savings as well as GHG emission reduction vs. measured baseline
@@
MPC Aachen, Germany five office spaces in a university building and a test hall at the E.ON Energy Research Center forced air and radiant systems with district heating for radiators and concrete core activation 14–33 minimize energy 5.3–22.2% energy savings vs. measured baseline
@@
MPC NY, US a subspace in a small commercial building with retail, office, and industrial spaces forced air system with five mini-split heat pumps and one single-stage, single-speed gas furnace 13 minimize cost and peak demand 27% cost and 23% peak demand vs. measured baseline
@@
MPC Sainte-Marthe-sur-le-Lac, QC, Canada six classrooms in a fully electric school building hybrid system with forced air in classrooms and radiant heating elsewhere. heat pumps, thermal storage, heat recovery ventilation. no cooling. $<1$ minimize cost 44% cost vs. measured baseline
@@
MPC Trondheim, Norway block A of a 6-story office building hydronic radiator heating system, ceiling-mounted ventilation diffusers, and automated sun-shading systems 40 minimize energy 10% energy savings vs. measured baseline
@@
RL Dübendorf, Switzerland one office in the HiLo Unit of the NEST Building, a modular innovation hub radiant hydronic system with TABS, hybrid VRF, and ventilation with heat recovery 49 minimize energy and improve comfort 68% reduction in temperature violations with similar energy consumption vs. simulated baseline
@@
MPC Incheon, South Korea single-zone office space in campus building at Inha University forced air system, direct expansion cooling, indoor and outdoor units, cooling-only 3 minimize energy and cost 25.2% energy and 33.7% cost vs. measured baseline
@@
Syracuse, NY, US an office in a small-to-medium commercial building at the Syracuse Center of Excellence forced air system with AHU, central VAV, and variable-speed fan 25–27 minimize energy 17.96–38.55% energy savings vs. measured baseline
@@
MPC Lausanne, Switzerland a multipurpose academic building, the Polydome, on the EPFL campus forced air system with AERMEC RTY-04 RTU and continuous ventilation 12 frequency regulation evaluation and minimize cost improved frequency regulation and 24.64–28.74% cost savings vs. measured baseline
80 [192]
2025
MPC Xi’an, China a research-focused office laboratory within a building forced air system with chiller, AHU, VAV, and variable-speed control 1 minimize energy 33.3–39.1% energy savings vs. measured baseline

### v6

@@ Appendix B Field demonstration summary tables
MPC Sargans, Switzerland three zones within an office building hydronic system with oil-fired boiler 46 minimize energy 31.9% energy savings vs. measured baseline
@@
MPC Bordeaux, France conference room in a university building radiant underfloor heating with AHU and fan coil unit for ventilation and backup heating 30 minimize energy 30–40% energy savings vs. simulated baseline
@@
MPC Long Beach, CA; Dayton, OH; NYC; Washington, DC, US office, courthouse, hospital, and high school buildings forced-air variable frequency drive systems 108–170 minimize energy -0.4–8.9% energy savings vs. measured baseline
@@
MPC West Lafayette, IN, US open-plan office space in a university research building hydronic radiant floor cooling, air-cooled chiller assumed 11 minimize energy 27% energy savings vs. simulated baseline
@@
MPC West Lafayette, IN, US and FL, US conference room in a university building and a retail store forced-air systems with packaged RTUs 3–60 minimize energy and peak demand 13–14% energy savings and 14.5–50% peak demand reduction vs. measured baseline
@@
MPC West Lafayette, IN, US three open-plan office spaces in a university research building radiant floor system in the first office, forced-air systems in the others 10 minimize energy and cost 7.8–64% energy savings and 34.7–78% cost savings vs. measured and simulated baseline
@@
MPC West Lafayette, IN, US open-plan office space in a university research building hydronic radiant floor cooling, air-cooled chiller assumed 19 improve comfort and minimize energy 23.1% thermal dissatisfaction reduction with energy increase vs. measured baseline
@@
RL Pittsburgh, PA, US single-level office building radiant heating system integrated with window mullions 78 minimize energy 16.7% energy savings vs. measured baseline
@@
MPC Singapore experimental cells in a building performance test facility hybrid system with active chilled beams and primary AHU-induced mixing 1 minimize energy 14.7–20% energy savings vs. measured baseline
@@
MPC Singapore single lecture room in a university building forced-air VAV with central plant and dedicated outdoor air system 14 minimize energy 16–20% energy savings vs. measured baseline
@@
MPC Varennes, QC, Canada single-story building with workstations and meeting rooms forced-air with electric boiler for heating and natural gas boilers as backup 70 minimize energy 22.2% energy savings vs. measured baseline
@@
MPC Hasselt, Belgium three-floor office building, underground garage and under-roof apartment excluded hybrid radiant and forced-air system with TABS and ground-source HP 35 minimize energy 53.5% energy savings vs. measured baseline
@@
MPC Singapore office and lecture theater in a multi-story institutional building forced-air with ceiling-mounted VAV and external fan coil unit 10 minimize energy 36.7–58.5% energy savings vs. measured baseline
@@
RL Pittsburgh, PA, US conference room in a university building forced-air VAV with variable-speed airflow 21 minimize energy 16.7% energy savings vs. measured baseline
@@
MPC Hamburg, Germany large government office building radiant heating with district hot water 90 minimize energy 30% energy savings vs. measured baseline
@@
MPC Brussels, Belgium five-floor building forced-air with central heating, cooling equipment, and AHUs 10–17 minimize energy 8.6–11.7% energy savings vs. simulated baseline
@@
MPC Singapore experimental cells in a building performance test facility hybrid system with active chilled beams and primary AHU-induced mixing 1–2 minimize energy 18.4–23.8% energy savings vs. measured baseline
@@
RL Berkeley, CA, US office room in a controlled test facility forced-air with variable-capacity AHU, shared chiller, PV, and battery 7 minimize cost 0.6–39.6% cost savings vs. measured baseline
@@
RL Seoul, South Korea climate chamber in controlled environment NR $<1$ improve comfort 10.9% predicted thermal discomfort (PPD) reduction vs. measured baseline
@@
MPC Heverlee, Belgium unoccupied test room in a research facility radiant heating via TABS, supplied by an electric boiler 57–69 compare MPC modeling approaches no model consistently outperformed others vs. measured data
@@
MPC West Lafayette, IN, US conference room in a university building forced-air with packaged RTUs 5–15 minimize cost 3.6–8.7% cost savings vs. measured baseline
@@
MPC Berkeley, CA, US two office floors with closed offices below and open offices above forced-air underfloor system with direct expansion cooling and heating 31 minimize energy 40% energy savings vs. measured baseline
@@
RL Nashville, TN, US three-story, energy-efficient mixed-use building forced-air with a dedicated outdoor air system, VRF, and dual AHUs 365 minimize energy 13.1–14.3% energy savings vs. measured baseline
@@
MPC Merced, CA, US central cooling plant at a university forced-air with chilled water, thermal storage, PV, and flexible chillers 7 maximize PV self-consumption, minimize peak demand and emissions 26.4% PV self-consumption increase, 9.8% peak demand reduction, and 9.6% emissions reduction vs. measured baseline
@@
MPC Blue Lake Rancheria, CA, US gas station convenience store forced-air RTUs with refrigeration, freezer, battery, and PV integration 365 minimize cost and peak demand 11.7% cost savings and 34% peak demand reduction vs. measured baseline
@@
MPC Campo Grande, Brazil operating and waiting rooms in a hospital forced-air cooling with AHUs and variable-speed chiller 4 minimize energy experimental savings unreported; simulated energy savings 4.76% vs. simulated baseline
@@
RL Singapore office in a net-zero university building hybrid system with dedicated outdoor air system and variable-speed ceiling fans 10 minimize energy 14% energy savings vs. measured baseline
@@
RL US university and mixed-use commercial buildings hybrid system with chiller plant, AHUs, fan coils, and central air distribution 90 minimize energy 9–13% energy savings vs. measured baseline
@@
MPC Singapore six-zone office space in a net-zero energy building forced-air with two fan coil units and individual zone VAVs 49 maximize PV self-consumption and building self-sufficiency 19.5% PV self-consumption increase and 10.6% building self-sufficiency increase vs. simulated baseline
@@
MPC West Lafayette, IN, US two identical adjacent private office spaces forced-air with individual VAV boxes and centralized AHU with reheat 1 minimize energy 28–35% energy savings vs. measured baseline
@@
MPC Zhejiang province, China three-floor airport terminal forced-air with air-cooled HPs, chilled water pumps, and AHUs 15 minimize energy 10.6–37.3% energy savings vs. measured baseline
@@
MPC Costa Mesa, CA, US six classrooms across two buildings in a K-12 school forced-air with six RTUs serving each classroom 30 minimize peak demand 22-–24% peak demand reduction vs. measured baseline
@@
MPC
Lausanne, Switzerland multipurpose academic building forced-air HP 4 minimize energy 18.4% energy savings vs. measured baseline
@@
MPC ON and QC, Canada central heating plants radiant steam system with non-condensing gas boilers for heating and hot water 48–67 minimize energy, cost, and GHG emissions 1.3–2.8% energy, cost, and GHG emissions reduction vs. measured baseline
@@
MPC Aachen, Germany five university office spaces and a test hall in an energy facility forced-air and radiant systems with district heating and concrete core activation 14–33 minimize energy 5.3–22.2% energy savings vs. measured baseline
@@
MPC NY, US subspace in a small mixed-use commercial building forced-air with mini-split HPs and a gas furnace 13 minimize cost and peak demand 27% cost savings and 23% peak demand reduction vs. measured baseline
@@
MPC Sainte-Marthe-sur-le-Lac, QC, Canada six classrooms in a fully electric school building hybrid with forced-air, radiant heating, HPs, and thermal storage $<1$ minimize cost and peak demand 44% cost savings and 47–95% peak demand reduction vs. measured baseline
@@
MPC Trondheim, Norway a block of a 6-story office building hydronic heating, mechanical ventilation, and sun-shading 40 minimize energy 10% energy savings vs. measured baseline
@@
RL Dübendorf, Switzerland office in a modular high-performance research building radiant hydronic TABS, hybrid VRF, and heat recovery ventilation 49 minimize energy and improve comfort 68% temperature violations reduction with similar energy consumption vs. simulated baseline
@@
MPC Incheon, South Korea single-zone office space in a university building forced-air direct expansion cooling system 3 minimize energy and cost 25.2% energy savings and 33.7% cost savings vs. measured baseline
@@
Syracuse, NY, US office in a small-to-medium university building forced-air AHU with VAV and variable-speed fan 25–27 minimize energy 18–38.6% energy savings vs. measured baseline
79 [192]
2025
MPC Xi’an, China university research laboratory forced-air with chiller, AHU, and VAV 1 minimize energy 33.3–39.1% energy savings vs. measured baseline
@@
MPC
Lausanne, Switzerland multipurpose academic building forced-air with rooftop unit and continuous ventilation 12 frequency regulation evaluation and minimize cost improved frequency regulation and 24.6–28.7% cost savings vs. measured baseline